The financial statements of Carla Vista Inc. are presented here: CARLA VISTA INC. Income Statement Year Ended December 31, 20$136,500 Liabilities and Shareholders Equity Current liabilities Accounts payable $89,000 Income tax payable 20,000 CurrentCalculate the following values and ratios for 2021. We provide the results for 2020 for comparative purposes. (Round current
Answer #1

Answer

1 Working capital

Working capital = current assets - current liabilities

2021 working capital = 401,900 - 136,500 = $265,400

2 Current ratio

Current ratio = current assets / current liabilities

Current ratio in 2021 = 401,900 / 136,500 = 2.94 : 1

3 Debt to total assets ratio

Debt to total assets = (total debt / total assets) × 100

Debt includes both long term liabilities and current liabilities.

Debt to total assets ratio

= ( 399,500 / 993,200 ) × 100 = 40.22%

4 Basic earning per share

= income available to common share holders / weighted average number of common shares outstanding

Basic earnings per share = 151,000 / 30,000 = $5.03

5 Price earning ratio

Price earning ratio = market price per share / Basic earnings per share

Price earning ratio

= 35 / 5.03 = 6.96 times

Working capital $265,400
Current ratio 2.94 : 1
Debt to total assets 40.22%
Basic earnings per share $5.03
Price earning ratio 6.96 times

ANALYSIS

1 Liquidity

Liquidity IMPROVED in 2021

Liquidity can be measured by using working capital and current ratio. The current ratio and working capital increase in 2021 compared to year 2020.

2 Solvency

Solvency DECLINED in 2021

The debt to total assets ratio is a measure of solvency. The debt to total assets ratio increased in 2021 compared to 2020. Increase in debt to assets ratio indicate the risk of solvency.

3 profitability

Profitability IMPROVED in 2021

Profitability can be measured by using basic earnings per share. The basic earning per share increased from 2020 to 2021. The increase in basic earnings per share indicates improvement in profitability. In 2020 basic earnings per share is $3.15 and in 2021 basic earnings per share is $6.96.

If YOU are like the answer please upvote.